# Improving Antibody Design with Force-Guided Sampling in Diffusion Models

### v1

@@ 3.2.1 Force Field
In the context of antibody design, the force field takes a protein $P$ (e.g., set of atom coordinates $x$) and computes the energy $U$. By calculating the gradient $\nabla U$, we can determine how $U$ varies with changes in atomic positions. This gradient indicates how to adjust each atom’s position to minimize the total energy of the protein structure. Lower energy configurations often correspond to more thermodynamically stable antigen-antibody complexes, which are typically associated with higher affinity and specificity (Ji et al., 2023). Using the relationship between energy and force, $-\nabla U(x)=F$, we can simulate the equations of motion to evolve this dynamical system according to the potential energy $U$.
@@ 3.2.2 DiffForce-$C_{\alpha}$: Interpolating Between $p_{\mathrm{data}}$ and $e^{-\kappa U(x_{0};C)}$
For simplicity, we consider the setting where the residues are fixed, and our focus is to guide the $C_{\alpha}$ atom coordinates with a prescribed force field. Rather than sample unconditionally from the data distribution, we are interested in sampling from the following combined distribution:
@@
where we use the notation $U(x_{0};C)$ to denote that $C$ is fixed throughout simulation. This induces a new distribution we wish to sample from that interpolates between the Boltzmann distribution $e^{-\kappa U(x_{0})}$  [^1] For brevity we have dropped the conditioning on $C$. and $p_{\mathrm{data}}(x_{0})$. One way to interpret this is to think of $p_{\mathrm{data}}(x_{0})$ as a prior and $e^{-\kappa U(x_{0})}$ acting as a likelihood of the form $p(y|x_{0})$. Thus $\pi_{0}(x_{0})$ is akin to a posterior of the form $p(x_{0}|y)$ that is in a way conditioned to make the binding energy small. However, unlike (Song et al., 2023; Komorowska et al., 2024), we do not have an explicit notion of the variable $y$ in this setting. We highlight that (Wang et al., 2024) concurrently explore an akin setting; however, their approach is focused on learning a new modified score while our work is focused on approximations during inference.
An alternate and akin approach is to construct $\pi_{0}$ as an annealed distribution (Neal, 2001) that is to form the weighted geometric mean $\pi_{0}\propto p_{\mathrm{data}}^{1-\beta}\exp(-\kappa U(x_{0}))^{\beta}$ for $\beta\in[0,1]$. This has the interpretation that we are now trying to sample from a distribution that is an interpolation between $p_{\mathrm{data}}(x_{0})$ and $e^{-\kappa U(x_{0})}$. By leveraging the weighted geometric mean of the distributions, we ensure that if one distribution suggests a particular outcome is extremely unlikely, it significantly influences the other, thus pulling the combined distribution towards more realistic outcomes. This method aligns well with our goal of generating high-quality samples with good binding energies, providing a balanced compromise between the two. In practice, however, we follow Equation 2 as it provides a form that is easier to tune and more in line with prior works on conditioning diffusion models. Due to this connection, we will refer to $\pi_{0}$ as the interpolating distribution. To sample from Equation 2 we must estimate the interpolating score $\nabla_{x_{t}}\pi_{t}(x_{t})$ (Chung et al., 2023):
@@
$=\pi_{0}(\mathbb{E}[x_{0}|x_{t}])\propto e^{-\kappa U(\mathbb{E}[x_{0}|x_{t}])}.$ (5)
Then, the approximate interpolating score is given by $\nabla_{x_{t}}\ln\pi_{t}(x_{t})\approx-\kappa\nabla_{x_{t}}U(\mathbb{E}[x_{0}|x_{t}]),$ and we can use Tweedie’s formula (Robbins, 1992) to compute $\mathbb{E}[x_{0}|x_{t}]$ given we have a good approximation of the score:
@@ 3.2.3 Force Guidance for Residue Types
The denoising process for amino acid orientations is captured through SO(3) elements, as described by (Leach et al., 2022) and implemented by (Luo et al., 2022, Equation 11), namely:
@@
where $H$ is a neural network that denoises the orientation matrix for amino acid $j$, $\mathcal{IG}_{\text{SO(3)}}$ denotes the isotropic Gaussian distribution on SO(3) parameterized by a mean rotation and a scalar variance, $\beta_{\tau}^{t}$ is the variance increase with the step $t$. To obtain the approximation $\hat{O}^{0}_{j}$ for amino acid orientation, we propose an approach of iteratively denoising the sample $O^{t}_{j}$ for $t$ iterations, where each iteration predicts the sample $O^{t-1}_{j}$. Namely, by iteratively applying Equation 11 until timestep $t$ reaches 0 for each amino acid $j$, we converge to the approximation $\hat{O}^{0}_{j}$, effectively reversing the forward diffusion process, more precisely:
@@ 3.2.4 Implementation
We derive a novel approach for guiding the sampling of $C_{\alpha}$ atom coordinates with a prescribed force. The algorithm can be summarized as follows:
@@
We introduce two hyperparameters, force scale ($\lambda_{sc}$) and force start ($\lambda_{st}$), to control the force’s impact during sampling. The force scale parameter dictates the magnitude of the force, gradually adjusted from $0.0$ to $\lambda_{sc}$ using a linear energy scheduling strategy. This parameter is applied to normalized forces as detailed in Appendix B. The force start parameter ($\lambda_{st}$) defines when force application begins, with a value of 0.3 indicating initiation of force at 70% of the sampling.
@@ Appendix D Structure Reconstruction
To calculate the energy, it is essential to reconstruct the full antibody-antigen complex $C$ along with the generated CDR region $R$. This process involves first reconstructing the complete 3D structure of the atoms in the CDR, following the pipeline outlined in (Luo et al., 2022). The reconstruction begins by determining the coordinates of the N, $C_{\alpha}$, C, O, and side-chain $C_{\beta}$ atoms, which are positioned relative to the $C_{\alpha}$ location and orientation of each amino acid (Engh and Huber, 2012). After these core atoms are reconstructed, the remaining side-chain atoms are built using the side-chain packing function in Rosetta (Alford et al., 2017). Once the CDR region is restored, the full antibody-antigen complex $C$ is reconstructed. With the complete structure (including the antibody with its 6 CDRs and framework, as well as the antigen), the energy of the complex can be calculated. This process is then iteratively performed for $\lambda_{st}\times 100$ timesteps, assuming diffusion occurs over $t=100$ timesteps. The iteration begins when forces are first applied at $\lambda_{st}$ and continues through the sampling process until the final timestep $t=0$.

### v2

@@ 3.2.1 Force Field
In the context of antibody design, the force field takes a protein $P$ (e.g., set of atom coordinates $x$) and computes the energy $U$. By calculating the gradient $\nabla U$, we can determine how $U$ varies with changes in atomic positions. This gradient indicates how to adjust each atom’s position to minimize the total energy of the protein structure. Lower energy configurations often correspond to more thermodynamically stable antigen-antibody complexes, which are associated with higher affinity (Ji et al., 2023). Using the relationship between energy and force, $-\nabla U(x)=F$, we can simulate the equations of motion to evolve this dynamical system according to the potential energy $U$.
@@ 3.2.2 DiffForce-$C_{\alpha}$ : Interpolating Between $p_{\mathrm{data}}$ and $e^{-\kappa U(x_{0};C)}$
For simplicity, we consider the setting where the residues are fixed, and our focus is to guide the $C_{\alpha}$ atom coordinates with a prescribed force field. Rather than sample unconditionally from the data distribution, we are interested in sampling from the following tilted distribution:
@@
where we use the notation $U(x_{0};C)$ to denote that $C$ is fixed throughout simulation. This induces a new distribution we wish to sample from that interpolates between the Boltzmann distribution $e^{-\kappa U(x_{0})}$  [^1] For brevity we have dropped the conditioning on $C$. and $p_{\mathrm{data}}(x_{0})$. One way to interpret this is to think of $p_{\mathrm{data}}(x_{0})$ as a prior and $e^{-\kappa U(x_{0})}$ acting as a likelihood of the form $p(y|x_{0})$. Thus $\pi_{0}(x_{0})$ is akin to a posterior of the form $p(x_{0}|y)$ that is in a way conditioned to make the binding energy small. However, unlike (Song et al., 2023; Komorowska et al., 2024), we do not have an explicit notion of the variable $y$ in this setting. We highlight that (Wang et al., 2024) concurrently explore an akin setting; however, their approach is focused on learning a new modified score while our is focused on approximations during inference.
An alternate and akin approach is to construct $\pi_{0}$ as a log-concave interpolation, as in annealed sampling (Neal, 2001), that is to form the weighted geometric mean $\pi_{0}\propto p_{\mathrm{data}}^{1-\beta}\exp(-\kappa U(x_{0}))^{\beta}$ for $\beta\in[0,1]$. This has the interpretation that we are now trying to sample from a distribution that is an interpolation between $p_{\mathrm{data}}(x_{0})$ and $e^{-\kappa U(x_{0})}$. By leveraging the weighted geometric mean of the distributions, we ensure that if one distribution suggests a particular outcome is extremely unlikely, it influences the other, thus pulling the combined distribution towards more realistic outcomes. This method aligns well with our goal of generating high-quality samples with good binding energies, providing a balanced compromise between the two. In practice, however, we follow Equation 2 as it provides a form that is easier to tune and more in line with prior works on conditioning diffusion models. Due to this connection, we will refer to $\pi_{0}$ as the interpolating distribution. To sample from Equation 2 we estimate the interpolating score $\nabla_{x_{t}}\ln\pi_{t}(x_{t})$ (Chung et al., 2023):
[…]
@@
$=e^{-\kappa U(\mathbb{E}[x_{0}|x_{t}];C)}.$ (7)
Then, the approximate interpolating score is given by $\nabla_{x_{t}}\ln\pi_{t}(x_{t})\approx-\kappa\nabla_{x_{t}}U(\mathbb{E}[x_{0}|x_{t}])+\nabla_{x_{t}}\ln p(x_{t}),$ and we can use Tweedie’s formula (Robbins, 1992) to compute $\mathbb{E}[x_{0}|x_{t}]$ given we have a good approximation of the score:
@@ 3.2.3 Force Guidance for Residue Types
The denoising process for amino acid orientations is captured via SO(3) elements, as described by (Leach et al., 2022) and implemented by (Luo et al., 2022, Equation 11):
@@
where $H$ is a neural network that denoises the orientation matrix for amino acid $j$, $\mathcal{IG}_{\text{SO(3)}}$ denotes the isotropic Gaussian distribution on SO(3) parameterized by a mean rotation and a scalar variance, $\beta_{\tau}^{t}$ is the variance increase with the step $t$. To obtain the approximation $\hat{O}^{0}_{j}$ for amino acid orientation, we propose an approach of iteratively denoising the sample $O^{t}_{j}$ for $t$ iterations, where each iteration predicts the sample $O^{t-1}_{j}$. Namely, by iteratively applying Equation 13 until timestep $t$ reaches 0 for each amino acid $j$, we converge to the approximation $\hat{O}^{0}_{j}$, effectively reversing the forward diffusion:
@@ 3.2.4 Implementation
We derive a novel approach for guiding the sampling of $C_{\alpha}$ atom coordinates with a prescribed force:
@@
We introduce two hyperparameters, force scale ($\lambda_{sc}$) and force start ($\lambda_{st}$). The $\lambda_{sc}$ parameter dictates the magnitude of the force, gradually adjusted from $0.0$ to $\lambda_{sc}$ using a linear scheduling strategy. This parameter is applied to normalized forces as detailed in Appendix B. The $\lambda_{st}$ parameter defines when force application begins, with a value of 0.3 indicating initiation of force at 70% of the sampling.
@@ Appendix D Structure Reconstruction
To calculate the energy, it is essential to reconstruct the full antibody-antigen complex $C$ along with the generated CDR region $R$. This process involves first reconstructing the complete 3D structure of the atoms in the CDR, following the pipeline outlined in (Luo et al., 2022). The reconstruction begins by determining the coordinates of the N, C, O, and side-chain $C_{\beta}$ atoms, which are positioned relative to the $C_{\alpha}$ location and orientation of each amino acid (Engh and Huber, 2012). After these core atoms are reconstructed, the remaining side-chain atoms are built using the side-chain packing function in Rosetta (Alford et al., 2017). Once the CDR region is restored, the full antibody-antigen complex $C$ is reconstructed. With the complete structure (including the antibody with its 6 CDRs and framework, as well as the antigen), the energy of the complex can be calculated. This process is then iteratively performed for $\lambda_{st}\times 100$ timesteps, assuming diffusion occurs over $t=100$ timesteps. The iteration begins when forces are first applied at $\lambda_{st}$ and continues through the sampling process until the final timestep $t=0$.